X-6y=11 and 2x-5y=1. Solve using substitution and solving for an ordered pair
Answer by stanbon(75887)

X-6y=11 and 2x-5y=1. Solve using substitution and solving for an ordered pair
---------------------
x = 6y+11
-----
Substitute and solve for "y":
2(7y+11) - 5y = 1
------
9y + 22 = 1
9y = -21
y = -7/3
============
Cheers,
Stan H.